OpenStudy (kymber):

Why doesn't Does 6÷2(1+2) equal 1?

OpenStudy (kymber):

Everywhere I look, people say it's 9. Shouldn't you follow PEMDAS? Parenthesis, 1+2 = 3 So then it's 6÷2(3) According to PEMDAS, you should do Multiplication and then Division, right? So shouldn't you do 2*3 = 6 And then 6÷6 = 1?

OpenStudy (anonymous):

no pemdas states that when multiplying and dividing you go from left to right. same for add and subtract. so its 6/2 AND THEN 3*3 = 9

OpenStudy (kymber):

Oh :/ I forgot the left to right part. Thanks!

OpenStudy (anonymous):

lol?